# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v1

@@ Appendix C Likelihood and priors
Appendix D Gaia Challenge logarithmic density slope profile recovery with an $\alpha\beta\gamma$ mass model
@@
Appendix E Mass estimators with non-Gaussian PDFs and multiple tracers
In this Appendix, we explore the behavior of the simple mass estimators using GravSphere2’s non-Gaussian PDFs (Appendix A) for both 100 tracers and a much larger number of tracer stars (10,000). The results are shown in Fig. 9, which compares the GravSphere2 recovery of the enclosed mass for the PlumCoreOM Gaia Challenge mock with both 100 and 10,000 tracers with the same for simple mass estimators. Notice that while biases improve following these PDFs, they still persist for the Wolf and Walker estimators for 100 tracers and (to a varying extent) in all of them for the 10,000 tracer case, while GravSphere2 remains unbiased.
Figure 9: Left: GravSphere2 recovery, using GravSphere2’s non-Gaussian PDFs (Appendix. A), of the enclosed mass profile for the PlumCoreOM mock galaxy with only LOS velocities for 100 tracers, as compared to the same for simple mass estimators (as marked). The bands show the 95% CL values. Right: Same as left but for 10,000 tracers. While non-Gaussian PDFs improve results, a degree of bias is present in all estimators, though to a lesser extent in the Errani estimator, where the bias becomes negligible (i.e. relative to statistical errors) for 100 tracers.
Appendix F Gaia Challenge binned profiles
Figs. 10 and 11 show binned projected velocity dispersion and kurtosis profiles for the Gaia Challenge mock sample.
Figure 10: Projected binned velocity dispersion and kurtosis profiles for the PlumCoreOM Gaia Challenge mock galaxy. LOS and PM velocity dispersion profiles are shown in the upper panels, whilst kurtosis profiles are shown in the bottom ones. The error bars denote 68 % CL regions from the median, whilst the colored bands denote the 95 % CL regions for 100 (green), 1,000 (blue), and 10,000 (gray) tracers. The LOS profile results are obtained from the LOS-only model, whilst the PM ones combine both data. Binned profiles are for the 10,000 tracer model in bins with an equal number of stars obtained by fitting the generalized PDFs (Eqs. (25) and (A)) for a given variance and kurtosis for each bin. The gray, dashed line denotes the projected half-light radius ($R_{1/2}$).
@@ Appendix F Gaia Challenge binned profiles
Appendix G Simulated dwarf galaxy binned profiles
@@
Figure 12: Projected binned projected velocity dispersion and kurtosis profiles for the Fornax-like simulated dwarf galaxy. The LOS and PM velocity dispersion profiles are shown in the upper panels, whilst kurtosis profiles are shown in the bottom ones. The error bars denote 68 % CL regions from the median, whilst the colored bands denote the 95 % CL regions for 100 (green), 1,000 (blue), and all tracers (gray). The LOS profile results are obtained from the LOS-only model, whilst the PM ones combine both data. Binned profiles are derived using all the bound and velocity-clipped tracers from the simulation in bins with of equal number fitting the generalized PDFs (Eqs. (25) and (A)) for a given variance and kurtosis for each bin. The gray, dashed line denotes the projected half-light radius ($R_{1/2}$).

### v2

@@ Acknowledgements.
Appendix A Gaia Challenge logarithmic density slope profile recovery with an $\alpha\beta\gamma$ mass model
@@
Appendix B Recovery comparison with priors
In Fig. 9 we show as an example how, even for our least constraining model of the Gaia Challenge sample – just LOS velocities of 100 tracers – the range of densities allowed by our priors (which is very similar across all the galaxies considered in our mock samples) spans several orders of magnitude and therefore has little effect on the result when using the data to constrain the mass model. We showed this by running a fit without any likelihood (just sampling the prior space) and plotting the range of densities spanned by the model in this case. Fig. 10 shows the recovery of the tracer surface density for the same mock galaxy using a binned profile with 10,000 photometric tracers. It is clear that the recovery is excellent and far in excess of the allowable range by the prior. In the cases we have considered, results will be essentially unchanged if one instead fixes the photometric profile to its best-fit value without additional marginalization. This can be used to accelerate run-times in GravSphere2.
Figure 9: Comparison of GravSphere2 recovery (95% CL from the median) with just 100 LOS tracers and the range spanned by the priors on the mass model for the PlumCoreOM mock galaxy from the Gaia Challenge suite (same as in Fig. 3).
Figure 10: Comparison of GravSphere2 recovery (95% CL from the median) with just 100 LOS tracers (10,000 for the binned photometric component) and the range spanned by the priors on the tracer surface density model for the PlumCoreOM mock galaxy from the Gaia Challenge suite (same as in Fig. 3).
Appendix C Mass estimators with non-Gaussian PDFs and multiple tracers
Here we explore the behavior of the simple mass estimators using GravSphere2’s non-Gaussian PDFs (Sec. 2.2) for both 100 tracers and a much larger number of tracer stars (10,000). The results are shown in Fig. 11, which compares the GravSphere2 recovery of the enclosed mass for the PlumCoreOM Gaia Challenge mock with both 100 and 10,000 tracers with the same for simple mass estimators. Notice that while biases improve following these PDFs, they still persist for the Wolf and Walker estimators for 100 tracers and (to a varying extent) in all of them for the 10,000 tracer case, while GravSphere2 remains unbiased.
Figure 11: Left: GravSphere2 recovery, using GravSphere2’s non-Gaussian PDFs (Sec. 2.2), of the enclosed mass profile for the PlumCoreOM mock galaxy with only LOS velocities for 100 tracers, as compared to the same for simple mass estimators (as marked). The bands show the 95% CL values. Right: Same as left but for 10,000 tracers. While non-Gaussian PDFs improve results, a degree of bias is present in all estimators, though to a lesser extent in the Errani estimator, where the bias becomes negligible (i.e. relative to statistical errors) for 100 tracers.
Appendix D Gaia Challenge binned profiles
Figs. 12 and 13 show binned projected velocity dispersion and kurtosis profiles for the Gaia Challenge mock sample.
Figure 12: Projected binned velocity dispersion and kurtosis profiles for the PlumCoreOM Gaia Challenge mock galaxy. LOS and PM velocity dispersion profiles are shown in the upper panels, whilst kurtosis profiles are shown in the bottom ones. The error bars denote 68 % CL regions from the median, whilst the colored bands denote the 95 % CL regions for 100 (green), 1,000 (blue), and 10,000 (gray) tracers. The LOS profile results are obtained from the LOS-only model, whilst the PM ones combine both data. Binned profiles are for the 10,000 tracer model in bins with an equal number of stars obtained by fitting the generalized PDFs (Eqs. (25) and (2.2)) for a given variance and kurtosis for each bin. The gray, dashed line denotes the projected half-light radius ($R_{1/2}$).
@@ Appendix D Gaia Challenge binned profiles
Appendix E Simulated dwarf galaxy binned profiles
@@
Figure 14: Projected binned projected velocity dispersion and kurtosis profiles for the Fornax-like simulated dwarf galaxy. The LOS and PM velocity dispersion profiles are shown in the upper panels, whilst kurtosis profiles are shown in the bottom ones. The error bars denote 68 % CL regions from the median, whilst the colored bands denote the 95 % CL regions for 100 (green), 1,000 (blue), and all tracers (gray). The LOS profile results are obtained from the LOS-only model, whilst the PM ones combine both data. Binned profiles are derived using all the bound and velocity-clipped tracers from the simulation in bins with of equal number fitting the generalized PDFs (Eqs. (25) and (2.2)) for a given variance and kurtosis for each bin. The gray, dashed line denotes the projected half-light radius ($R_{1/2}$).